OpenStudy (anonymous):

Use the quadratic formula to solve: 2y^2 – 3y = 2.

OpenStudy (anonymous):

2y^2 - 3y - 2 = 0 is the form that conforms to: ay^2 + by + c = 0 so, a=2, b=-3, c=-2 and: \[y = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\] So, just substitute and simplify.

OpenStudy (luigi0210):

She's the type that likes to wait for the answer instead of doing it themselves :l

OpenStudy (anonymous):

I don't know how to do it exactly

OpenStudy (luigi0210):

I'm sorry.. but I've done questions with you and I've pretty much had to do them all for you :/

OpenStudy (luigi0210):

You could at least give some effort in trying to understand :/

OpenStudy (anonymous):

It's all there. All the prep work, theory, and explanation. At this point, it's up to you, @desiraedallas1999 . Otherwise, we end up giving away answers and that's not the Openstudy way.

OpenStudy (luigi0210):

I'm sorry if that sounded mean..

OpenStudy (anonymous):

Where you see an "a", put a "2" in. Where you see an "b", put a "-3" in. Where you see an "c", put a "-2" in.

OpenStudy (anonymous):

y=-3+or-/-3^2-4(2)(2)/(2)(2)

OpenStudy (anonymous):

Watch that sign on that first -3, it's supposed to be: -(-3) = 3

OpenStudy (anonymous):

And "c" is -2, not 2.

OpenStudy (anonymous):

now I am down to 3/7/4